# Exploiting Euclidean Distance Field Properties for Fast and Safe 3D planning with a modified Lazy Theta*

### v1

@@ 4.1 Case study
Figure 5: Selection of neighbours in 2D. The current node is in the coordinates’ center, $O_{1}$ and $O_{2}$ are the obstacles, and $G$ the goal. The grid in the centre shows the eight neighbours of the current node. First case (a): $G$ rotates with a small radius around $O_{1}$ (dotted line in red), and $O_{1}$ rotates around the current node (dashed line in red). Second case (b): $G$ rotates around $O_{2}$ (dotted line in blue) with a large radius, and $O_{2}$ rotates around the current node (dashed line in blue). The difference between both setups is that in the second one the current node might fall between $G$ and $O_{2}$.
@@ 5 Experimental results
This section presents experimental results in different 3D scenarios in order to analyze the robustness and performance of the FS-Planner. The method is benchmarked by considering different number of the visibility neighbours selected among all the nodes. The FS-Planner is tested with different numbers of visibility neighbours in 3D: $9,10,11,13,15$ or $17$ nodes. The proportion of nodes selected relative to all neighbours is similar to the study in 2D ($3$ or $5$ out of $8$), that is, from $9$ to $17$ out of $26$. In all the cases, the neighbours are selected around the best candidate, $\hat{u}$, (see Section 4) except when $10$ nodes are selected. In this particular case, we consider $9$ nodes around the best candidate, $\hat{u}$, and the node opposite to it. A third method is implemented depending on the angle between the minimum gradient unit vector, the unit vectors of $s_{edf}$, and the goal unit vector, the unit vector of $s_{goal}$. If the angle is smaller than $90^{o}$ we consider a number of neighbours and if the angle is greater than $90^{o}$ a larger number of neighbours is considered. For example, $9-11$ considers $9$ neighbours when the angle is smaller than $90^{o}$.
@@
We will benchmark the FS-Planner against the CALT*+M1 planner presented in [13], and the A* algorithms. All planners compute the path on the same EDF previously generated, so the computation times are exclusively referred to the generation of the path. Both FS-Planner and A* make use of the same EDF-based cost function (2). As A*+M1 in [13] mentioned in Section 1, CALT*+M1 considers an EDF as environment representation and a cost function that integrates costs related to the distance to obstacles. The results in [13] showed that the CALT*+M1 provided similar computation times to A*, while smoothness and safety are clearly improved. Moreover, A* algorithm is used as baseline in order to compare them considering the three algorithms the same cost function (2).
The scenarios considered are presented in Fig. 7: S1, like letter H, does not consider segments ”a” and ”c”, and the path goes from 1 to 3 point (see Fig. 7-left); S2, like letter U inverted, it does not consider segments ”b” and ”c” and the path goes from point 2 to 4 (see Fig. 7-left); S3 does not consider segment ”b”, and the path goes from point 2 to 4 (see Fig. 7-left); S4 is the maze shown in Fig. 7-right and the path goes from point 5 to 6; S5 is the same building considered in [13] (see Fig. 9). The objective is to test the FS-Planner in challenging scenarios, with narrow openings and windows considering convex and non-convex obstacles.
@@
Figure 7: Scenarios considered in the tests performed.
@@
The computer used to perform the tests are: Intel Core i7-10700 CPU 2.90GHz, 32 GB RAM, with Ubuntu OS 20.04 LTS. All the planning algorithms have been developed in C++ language and has been integrated into ROS (Robot Operating System) Noetic distribution.
@@ 5.1 Dependency on the number of visibility neighbours
This section shows the performance of the FS-Planner depending on the number of visibility neighbours and how it always outperforms CALT*+M1. We focus on the computation time and the explored nodes.
@@
Figure 9 shows how the selection of visibility neighbours reduces the explored nodes with the FS-Planner regarding CALT*+M1. The path goes from one building facade to another turning in a corner (the first path considered in scenario S5). Note that the building is not explored inside with the FS Planner when considering the gradient information.
Figure 9: Explored nodes in red to compute a path that goes from one building facade to another turning in a corner with: (left) CALT*+M1 and (right) FS-Planner considering 11 visibility neighbours.
@@ 5.2 Random tests
This section presents an evaluation with random tests to compare the CALT*+M1 and the FS-Planner using the A* algorithm as baseline. All planners make use of a previously computed EDF. Both FS-Planner and A* use the same cost function (2). Note that we are interested in the computation times of each planner on EDF and not in the EDF generation. The metrics to evaluate the efficiency, safety and smoothness are as in [13] (see Table 1): path length (length), clearance of obstacles (mean distance related to safety), computation time (time related to efficiency) and heading changes of the path (mean angle related to smoothness). We run twenty path planning tests with different start and goal positions distributed randomly in each scenario. Twenty simulations are performed in each problem to calculate the mean computation time. It is considered $c_{w}=500$ and $LoS=1$ in all the cases. Note that the difference between the bounds of (7) is small ($0.5$), so using the upper bound in (11 provides a good approximation for both convex and non-convex shapes. As a result, we do not need to infer the type of obstacle. The number of visibility neighbours are: 9-11, 10, 11 and 11-13. The solution is always found in the tests shown in this section, but the implementation of the FS-Planner also considers the case that no solution is found. In that case the FS-Planner will be run again considering $17$ neighbours to ensure the solution.
Table 1 summarises the core results in each scenario by using the A* algorithm as baseline, considering (2), in order to compare them. The ratio is calculated as $\frac{algorithm}{A*}$, and the results show the mean of the twenty tests performed and the standard error of the mean which represents the uncertainty about the mean. The best values are highlighted in bold.
@@
All the results show that the time rates of the FS-Planner considering different number of visibility neighbours are lower than the ones of CALT*+M1, as in Fig. 8. Even the FS-Planner shows times lower than A* algorithm over EDF using (2) in scenarios S5 and S6 (see Table 1). Moreover, the explored nodes with CALT*+M1 and the FS-Planner are always lower than the explored nodes with A* algorithm except in S4 that CALT*+M1 is slightly greater (see Table 1-S4). The length is usually very close and similar to the one of A* with CALT*+M1 and FS-Planner. It is greater than the length of A* only with the FS-Planner considering $9-11$ neighbours in S3. This scenario is really challenging and the computed path is longer inside of the U almost closed.
The FS-Planner presents the best values of safety in all scenarios with respect to CALT*+M1 except in S4 in which they are very similar. Also it is safer than A* in scenarios S1, S2, S3 and S6, and very much alike in scenarios S4 and S5. Regarding the smoothness, the FS-Planner and CALT*+M1 compute similar and much smoother paths than A* algorithm, but not in S3 with $9-11$ and $11$ neighbours. Figure 10-left depicts the path computed by A* and the FS-Planner in S4 to show how the smoothness is improved.
@@
It is important to highlight the results obtained in the real scenario considered in [13]. We consider two paths in this real scenario: the first one goes from one building facade to another turning in a corner (see Table 1-S5 and Fig. 10-right); and the second one passes through the window from the outside to the inside (see Table 1-S6). The FS-Planner shows better values than CALT*+M1. Moreover, the computation times of the FS-Planner with 9-11 and 11 neighbours are lower than the ones with A* algorithm. The length and safety are very similar and the paths are much smoother than the ones computed by A* algorithm. Figure 10-right illustrates the improvement of the smoothness with the FS-Planner in S5. Note that the path computed by FS-Planner 9-11 is also smoother than the one computed by A* on the Z axis.
@@
The results demonstrate that the FS-Planner, which exploits the EDF properties and reduces the number of visibility neighbours, works better than CALT*+M1 in terms of computation times, explored nodes, safety and smoothness. Although the length is very similar in both planners, CALT*+M1 computes slightly shorter paths. Regarding A* algorithm over EDF using the cost function (2), the smoothness of the path computed by FS-Planner is considerably improved and the times too in some scenarios. Therefore, we can conclude the FS-Planner improves both CALT*+M1 and A* algorithm over EDF using cost function (2). Moreover, both the FS-Planner and CALT*+M1 do not make use of post-processing methods to achieve smoother paths.
@@
Figure 10: Paths considered in S4 (left) and S5 (right): A* in green and FS-Planner 9-11 in dark green.
@@
Table 1: Evaluation with metrics in the scenarios considered for CALT*+M1 and the FS-Planner considering 9-11, 10, 11, 11-13 and 13 neighbours with $LoS=1$ relative to A*: T is the ratio of the computation time, L is the ratio of the length, N is the ratio of the explored nodes, MD is the ratio of the mean distance to the closer obstacles and MA is the mean angle.
@@
S1 T $\downarrow$ $3.653\pm 0.073$ 1.666$\pm$0.053 $2.882\pm 0.046$ $2.297\pm 0.062$ $2.811\pm 0.052$
L $\downarrow$ 0.966$\pm$0.004 $1.020\pm 0.022$ $1.011\pm 0.009$ $1.001\pm 0.013$ $0.990\pm 0.011$
N $\downarrow$ $0.988\pm 0.001$ 0.541$\pm$0.009 $0.9824\pm 0.005$ $0.733\pm 0.014$ $0.867\pm 0.004$
MD $\uparrow$ $0.958\pm 0.017$ $0.884\pm 0.115$ 1.158$\pm$0.119 $0.992\pm 0.071$ $0.943\pm 0.056$
MA $\downarrow$ $0.471\pm 0.030$ $0.532\pm 0.046$ $0.467\pm 0.022$ $0.469\pm 0.049$ 0.443$\pm$0.035
S2 T $\downarrow$ $3.461\pm 0.020$ 1.827$\pm$0.017 $2.830\pm 0.055$ $2.178\pm 0.021$ $2.660\pm 0.025$
L $\downarrow$ 0.974$\pm$0.002 $1.033\pm 0.006$ $0.987\pm 0.006$ $1.019\pm 0.005$ $1.005\pm 0.005$
N $\downarrow$ $0.998\pm 0.001$ 0.634$\pm$0.003 $0.981\pm 0.001$ $0.761\pm 0.002$ $0.877\pm 0.004$
MD $\uparrow$ $1.020\pm 0.038$ 1.050$\pm$0.061 $0.951\pm 0.034$ $0.978\pm 0.049$ $1.008\pm 0.058$
MA $\downarrow$ $0.524\pm 0.014$ $0.576\pm 0.025$ $0.508\pm 0.011$ 0.540$\pm$0.016 $0.459\pm 0.013$
S3 T $\downarrow$ $3.398\pm 0.041$ 1.819$\pm$0.018 $2.660\pm 0.0174$ $2.152\pm 0.016$ $2.619\pm 0.022$
L $\downarrow$ 0.917$\pm$0.001 $1.351\pm 0.012$ $0.980\pm 0.005$ $1.075\pm 0.008$ $1.036\pm 0.007$
N $\downarrow$ $0.998\pm 0.001$ 0.642$\pm$0.004 $0.977\pm 0.002$ $0.779\pm 0.005$ $0.874\pm 0.005$
MD $\uparrow$ $0.971\pm 0.031$ $0.678\pm 0.017$ 1.042$\pm$0.030 $0.796\pm 0.021$ $0.934\pm 0.047$
MA $\downarrow$ 0.530$\pm$0.007 $1.661\pm 0.046$ $0.548\pm 0.021$ $1.099\pm 0.048$ $0.638\pm 0.031$
S4 T $\downarrow$ $3.566\pm 0.020$ 1.328$\pm$0.018 $2.732\pm 0.019$ $2.100\pm 0.017$ $2.547\pm 0.016$
L $\downarrow$ 0.988$\pm$0.004 $0.998\pm 0.006$ $0.989\pm 0.005$ $0.996\pm 0.003$ $0.995\pm 0.002$
N $\downarrow$ $1.004\pm 0.001$ 0.463$\pm$0.004 $0.969\pm 0.002$ $0.738\pm 0.003$ $0.833\pm 0.004$
MD $\uparrow$ 0.991$\pm$0.004 $0.974\pm 0.010$ $0.972\pm 0.010$ $0.969\pm 0.004$ $0.989\pm 0.003$
MA $\downarrow$ $0.588\pm 0.015$ 0.537$\pm$0.007 $0.545\pm 0.018$ $0.566\pm 0.011$ $0.556\pm 0.010$
S5 T $\downarrow$ $2.577\pm 0.033$ 0.782$\pm$0.034 $1.791\pm 0.052$ $0.790\pm 0.034$ $1.699\pm 0.028$
L $\downarrow$ $0.969\pm 0.003$ $0.967\pm 0.009$ $0.962\pm 0.005$ 0.952$\pm$0.004 $0.968\pm 0.004$
N $\downarrow$ $0.997\pm 0.004$ 0.385$\pm$0.014 $0.952\pm 0.005$ $0.406\pm 0.013$ $0.760\pm 0.004$
MD $\uparrow$ $0.989\pm 0.006$ $0.993\pm 0.014$ $0.964\pm 0.005$ $0.973\pm 0.007$ 0.995$\pm$0.007
MA $\downarrow$ $0.430\pm 0.013$ $0.445\pm 0.012$ $0.417\pm 0.009$ $0.414\pm 0.018$ 0.413$\pm$0.016
S6 T $\downarrow$ $2.807\pm 0.029$ 0.789$\pm$0.063 $2.076\pm 0.060$ $0.862\pm 0.119$ $1.722\pm 0.037$
L $\downarrow$ 0.972$\pm$0.043 $1.062\pm 0.027$ $0.996\pm 0.015$ $1.028\pm 0.036$ $0.976\pm 0.011$
N $\downarrow$ $0.999\pm 0.003$ 0.387$\pm$0.033 $0.974\pm 0.026$ $0.427\pm 0.066$ $0.727\pm 0.012$
MD $\uparrow$ $0.987\pm 0.007$ $0.926\pm 0.021$ 1.018$\pm$0.019 $0.948\pm 0.013$ $0.973\pm 0.012$
MA $\downarrow$ $0.670\pm 0.052$ $0.730\pm 0.063$ $0.588\pm 0.037$ $0.624\pm 0.044$ 0.571$\pm$0.033

### v2

@@ 4.1 Case study
Figure 5: Selection of neighbours in 2D. The current node, $S$, is in the coordinates’ center, $O_{1}$ and $O_{2}$ are the obstacles, and $G$ the goal. The grid in the centre shows the eight neighbours of the current node. As an analogy with the Sun–Earth–Moon system, in each case $S$ can be viewed as the Sun, $O_{1}$ and $O_{2}$ as the Earth, and $G$ as the Moon. Thus, points $O_{1}$ and $O_{2}$ rotate around $S$ like the Earth around the Sun, while G rotate around O1 and O2 in each case, respectively, like the Moon around the Earth. First case (a): $G$ rotates with a small radius around $O_{1}$ (dotted line in red). Second case (b): $G$ rotates around $O_{2}$ (dotted line in blue) with a large radius. The difference between both setups is that in the second one the current node might fall between $G$ and $O_{2}$.
@@ 5 Experimental results
This section presents experimental results in different 3D scenarios in order to analyze the robustness and performance of the FS-Planner. The method is benchmarked by considering different number of the visibility neighbours selected among all the nodes. The FS-Planner is tested with different numbers of visibility neighbours in 3D: $9,10,11,13,15$ or $17$ nodes. The proportion of nodes selected relative to all neighbours is similar to the study in 2D ($3$ or $5$ out of $8$), that is, from $9$ to $17$ out of $26$. In all the cases, the neighbours are selected around the best candidate, $\hat{u}$, (see Section 4) except when $10$ nodes are selected. In this particular case, we consider $9$ nodes around the best candidate, $\hat{u}$, and the node opposite to it. A third method is implemented depending on the angle between the minimum gradient unit vector, the unit vectors of $s_{edf}$, and the goal unit vector, the unit vector of $s_{goal}$. If the angle is smaller than $90^{o}$ we consider a number of neighbours and if the angle is greater than $90^{o}$ a larger number of neighbours is considered. For example, 9119-11 considers $9$ neighbours when the angle is smaller than $90^{o}$.
@@
We will benchmark the FS-Planner against the CALT*+M1 planner presented in [2], and the A* algorithms. All planners compute the path on the same EDF previously generated, so the computation times are exclusively referred to the generation of the path. Both FS-Planner and A* make use of the same EDF-based cost function (2). As A*+M1 in [2] mentioned in Section 1, CALT*+M1 considers an EDF as environment representation and a cost function that integrates costs related to the distance to obstacles. The results in [2] showed that the CALT*+M1 provided similar computation times to A*, while safety and heading changes between consecutive segments of the path are clearly improved. Moreover, A* algorithm is used as baseline in order to compare them considering the three algorithms the same cost function (2).
The simulated scenarios generated to test the planners are presented in Fig. 7: S1, like letter H, (see Fig. 7(a)); S2, like letter U inverted, (see Fig. 7(b)); S3 (see Fig. 7(c)); S4 is the maze shown in Fig. 7(d)). The tests have also been carried out in more realistic environments and are presented in Fig. 8: S5 and S6 consider the same building used in [2] ( see Fig. 8(a), 10 and 12). Figure 8 also shows the real-world outdoor environments used to test the planners; S7 considers point clouds derived from real-world data collected in an outdoor environment on the Nanyang Technological University (NTU) campus, referred to as the EEE environment in [18] (see Fig. 8(b)); S8 considers the point clouds derived from real data in the Computer Science campus at University of Freiburg [^3] http://ais.informatik.uni-freiburg.de/projects/datasets/octomap/ (see Fig. 8(c)). The objective is to test the FS-Planner in challenging scenarios, with buildings, narrow openings, trees and windows considering convex and non-convex obstacles. Note that we consider two paths in the scenario of [2] (see Fig. 8(a)): the first one goes from one building facade to another turning in a corner defined by S5 and E5 points (see Fig. 8(a), and Fig. 12); and the second one passes through the window from the outside to the inside defined by S6 and E6 poinsts (see Fig. 8(a)). Regarding S7 and S8, it is important to note that the start (S) and end (E) points are located close to ground level and below the tree canopies. This aspect can be more clearly appreciated in the corresponding path visualizations shown in Fig. 13 and 14
@@
Figure 7: Scenarios considered in the tests performed: (a) S1, (b) S2, (c) S3 and (d) S4. S is the start point and E the end point of the paths considered in each scenario.
Figure 8: Real-world outdoor environments considered in the tests performed: (a) S5 and S6, (b) S7 and (c) S8. S is the start point and E the end point of the paths considered in each scenario.
@@
The computer used to perform the tests is: Intel Core i7-10700 CPU 2.90GHz, 32 GB RAM, with Ubuntu OS 20.04 LTS. All the planning algorithms have been developed in C++ language and has been integrated into ROS (Robot Operating System) Noetic distribution.
@@ 5.1 Dependency on the number of visibility neighbours
This section shows the performance of the FS-Planner depending on the number of visibility neighbours and how it always outperforms CALT*+M1. We focus on the computation time and the explored nodes. We evaluate different numbers of selected neighbours to characterize the behavior of the FS-Planner. Since the FS-Planner modifies the original Lazy Theta* through the introduction of a gradient-based neighbour-selection mechanism, it is important to show how this mechanism behaves as the number of visibility neighbours changes.
@@
Figure 10 illustrates how the gradient-based visibility neighbour selection significantly reduces the number of explored nodes in the FS-Planner compared to CALT*+M1. The path considered corresponds to the first path in scenario S5, where the trajectory moves from one building façade to another while turning around a corner (see Figure 11-(right)). Note that CALT*+M1 explores a portion of the building’s interior, whereas the FS-Planner, by incorporating gradient information, avoids this exploration and focuses the search on the relevant region of the environment.
Figure 10: Explored nodes in red to compute a path that goes from one building facade to another turning in a corner with: (a) CALT*+M1 and (b) FS-Planner considering 11 visibility neighbours.
@@ 5.2 Random tests
This section presents an evaluation with random tests to compare the CALT*+M1 and the FS-Planner using the A* algorithm as baseline. All planners make use of a previously computed EDF. Both FS-Planner and A* use the same cost function (2). Note that we are interested in the computation times of each planner on EDF and not in the EDF generation. The metrics used to evaluate efficiency, safety, and smoothness follow the definitions in [2] (see Table 1): path length, computation time and number of explored nodes (related to efficiency), clearance to obstacles (related to safety), and a metric of geometric smoothness based on heading changes between consecutive path segments. For each scenario, we perform twenty planning tests with randomly distributed start and goal positions. Twenty simulations are performed in each problem to calculate the mean computation time. It is considered $c_{w}=500$ and $LoS=1$ in all the cases. Note that the difference between the bounds of (7) is small ($0.5$), so using the upper bound in (11 provides a good approximation for both convex and non-convex shapes. As a result, we do not need to infer the type of obstacle. The number of visibility neighbours are: 9-11, 10, 11 and 11-13. The solution is always found in the tests shown in this section, but the implementation of the FS-Planner also considers the case that no solution is found. In that case the FS-Planner will be run again considering $17$ neighbours to ensure the solution.
Table 1 summarises the main results for each scenario using A* algorithm, considering (2), as baseline for comparison. For each metric, the ratio is calculated as $\frac{algorithm}{A*}$, allowing a direct assessment of the relative performance with respect to the baseline. Table 1 reports the mean over twenty runs together with the standard error of the mean, which quantifies the uncertainty of the estimated average. Lower values of T, L, N and MA indicate better performance, whereas higher values of MD denote safer paths. The best values are highlighted in bold.
@@
All the results show that the computation times (T) and number of explored nodes (N) of the FS-Planner considering different number of visibility neighbours are lower than the ones of CALT*+M1, as in Fig. 9. Particularly, the FS-Planner configurations with 9–11 visibility neighbours obtain the lowest computation times and the smallest number of explored nodes in all eight scenarios. Even the FS-Planner shows times lower than A* algorithm over EDF using (2) in scenarios S5 and S6 (see Table 1). The length is usually very close and similar to the one of A* with CALT*+M1 and FS-Planner. It is greater than the length of A* only with the FS-Planner in some cases in S3 and S8. S3 is really challenging and the computed path is longer inside of the U almost closed.
Regarding safety (MD) and geometric smoothness (MA), the FS-Planner obtains the best values in six out of the eight scenarios. In the scenario S4 and S8, its performance is very close to that of CALT*+M1, showing that reducing the number of explored neighbours does not degrade the safety and geometric smoothness of the computed path. Also it is safer than A* in scenarios S1, S2, S3, S6 and S7, and very much alike in scenarios S4, S5 and S8. Figure 11 depicts the paths computed by A* and the FS-Planner in scenario S4. Both paths are safe, but the one generated by the FS-Planner is geometrically smoother, particularly in turns and during ascending or descending segments as is shown in Table 1.
@@
It is important to highlight the results obtained in the real scenarios considered in S5, S6, S7 and S8. The FS-Planner shows better values than CALT*+M1. Moreover, the computation times of the FS-Planner with 9-11 and 11 neighbours are lower than the ones with A* algorithm. The length and safety are very similar and the paths are much geometrically smoother than the ones computed by A* algorithm. Figure 12 illustrates the improvement of the geometric smoothness with the FS-Planner in S5. Note that the path computed by FS-Planner 9-11 is geometrically smoother than the one computed by A* during ascending segments (see Figure 12-right). Regarding real-world scenarios, S7 and S8, Figures 13 and 14 also illustrate the improvement of the smoothness with the FS-Planner.
@@
The results demonstrate that the FS-Planner, which exploits the EDF properties and reduces the number of visibility neighbours, works better than CALT*+M1 in terms of computation times, explored nodes, safety and geometric smoothness. Although the length is very similar in both planners, CALT*+M1 computes slightly shorter paths. Regarding A* algorithm over EDF using the cost function (2), the geometric smoothness of the path computed by FS-Planner is considerably improved and the times too in some scenarios. Therefore, we can conclude the FS-Planner improves both CALT*+M1 and A* algorithm over EDF using cost function (2). Moreover, both the FS-Planner and CALT*+M1 do not make use of post-processing methods to achieve geometrically smoother paths.
On the other hand, analyzing different numbers of selected neighbours allows us to empirically assess the robustness and scalability of the FS-Planner, and to clarify how the number of selected neighbours influences the main performance metrics (computation time, explored nodes, safety, and geometric smoothness). This evaluation confirms that the FS-Planner maintains high performance even with reduced neighbour sets.
Finally, it also is important to note that the experimental results provide empirical evidence of the behaviour of the FS-Planner, but they do not constitute formal guarantees of completeness or optimality. Although the fallback mechanism, where the FS-Planner is executed again considering all the neighbours, was never activated in any of the tested scenarios, including complex 3D and real-world environments, this observation should not be interpreted as a proof of completeness. Instead, it indicates that the neighbour selection policy behaves robustly across all evaluated scenarios.
Moreover, the twenty random planning tests performed per scenario indicate that the main properties and behaviour of the FS-Planner are consistently preserved across different configurations and situations. Although this does not constitute statistically exhaustive evidence, we believe that the results are very positive in terms of the quality of the generated paths and the robustness of the FS-Planner, given the wide variety of challenging scenarios considered.
@@
Figure 11: Paths considered in S4 in 2D (left) and 3D (right): A* in purple and FS-Planner 9-11 in black.
Figure 12: Paths considered in S5 in 2D (left) and 3D (right): A* in purple and FS-Planner 9-11 in black.
Figure 13: Paths considered in S7 in 2D (left) and 3D (right): A* in purple and FS-Planner 9-11 in black.
Figure 14: Paths considered in S8 in 2D (left) and 3D (right): A* in purple and FS-Planner 9-11 in black.
@@
Table 1: Evaluation with metrics in the scenarios considered for CALT*+M1 and the FS-Planner considering 9-11, 10, 11, 11-13 and 13 neighbours with $LoS=1$ relative to A*: T is the ratio of the computation time (related to efficiency), L is the ratio of the path length, N is the ratio of the explored nodes, MD is the ratio of the mean distance to the closest obstacles (related to safety) and MA is the mean angle (related to geometric smoothness).
@@
[…]